if each of base angles of an isoceles triangle is 5 less than twice the vertix angle. find the number of each angle?
Answer by mananth(16946)

let vertex angle = x
the base angles will be 2x-5
..
2x-5+2x-5+x = 180 ( sum of all angles = 180 deg)
5x-10=180
5x=180+10
5x=190
x=190/5
x= 38 vertex
2x-5 is the base angle = 2*38-5
=76-5
=71 deg.
The angles are 38,71,71